LBC Verification and Checkout after Summer Shutdown 2011

Work done over the Summer 2011

  • Red Chip Number 3 re-installed
    • Update LBC Description to show correct data (quantum efficiency, chip ID) for this specific chip (OPK to do)
  • New "C Type" CCD controllers on Blue
  • New software driver
  • Red shutter timing problem fixed; "due to inner electronics of the controller and not a mechanical problem" Fernando wrote.
  • dofpia updated to write logs to /home/LBCeng/FPIAlogs/ in the form (blue/red)YYYY-MM-DD.Log and (blue/red)YYYY-MM-DD.TecLog
  • Computers moved from treehouse to Computer Room by Control Room (Computer Room A)

Current status of LBC (2 September 2011)

  • Both LBCs show higher than usual thermal link temperatures, >200K for both channels. At the end of June/beginning of July these values were of order <~200 K for Blue and ~165 K for Red. Now Blue is stable at ~210 K and Red varying but reaching a minimum at ~190 K.
  • For Red Camera, thermal link temperature was very variable, and changed with telescope orientation. Temperature at detector correlated with that at thermal link, see plot below. Decided to dismount Red Camera for Fernando to look at. On schedule to be pumped, cooled and back on telescope for Saturday night.
  • 4 September: LBC-Red cool but not yet mounted. All day Fernando & Andrea worked on the tftp (data archiving) problem. LBC-Blue is working now.

There are some new items inside the control sw that we would like to test during the technical nights, and we think my help in case of bugs 
could be precious to all. During the tests i did with Muaro last days the system looked up more robust than before and a lot of OBs ran smoothly, 
of course there was no interaction with TCS so we consider those only a partial test.
The user interface is unchanged but now you are working with new sw drivers and new (C Type) ccd controller on the BLUE. The Read Out Time 
of the Blue despite of the new controller is almost the same because is not possible to speed up the vertical transfer and the detector wipeing due 
to the thin silicon higher capacitance (vs the red CCDs) that doesn't allow a faster transfer.
The shutter problem is fixed and it was due to the inner electronic of the controller not a real mechanical problem.

Test dofpia version.

  • Log file should be writable by all: Start dofpia as LBTO, stop, start as okuhn, stop, start as INAF, stop, start, ... After the file is created, start a new xterm and run
tail -f /home/LBCeng/FPIAlogs/filename.Log
  • Does the file update? Does dofpia crashes (crashes because of not finding pupils don't count here).
    • If so, OPK to check this version in to svn.

Summary of Standard LBC Checkout Tasks

Daily:
  • shut down UI and reboot the CMU before observing starts (I don't always do this, but it is recommended because of a shared memory leak. Run top -OPK).
  • check CMU system time vs our GPS: from CMU cmd line execute: "ntpdate -q 10.144.0.211"

Check stability of cameras and data flow:
  • Obtain biases and a dark. OBs for this are in the Calib_OBs package: Calib_OBs/BIASDARK and are 25Bias_Bino.ob and 2Darks_Bino.ob.
  • Check these. (can start up RB_Science script to view these as they are taken) Note the median levels for each chip. The most current values are at Detector Characteristics

Flat Fields:
We need to determine the exposure time scalings which produce similar counts in Blue and Red images.
  • These OBs do not have target information, they will observe at the "Current" position. So before running these, ask the TO to slew to track at the standard flat field position: RA=LST-00:30 and DEC=27 degrees.
  • There are two pairs of OBs which will generate a set of flats through all filters from which we can measure relative count rates. All red filters are measured against V-BESSEL and all blue filters against r-SLOAN (red). These OBs are in /home/LBTO/Calib_OBs/SKYFLAT_Scalings/ are are called:
    • Blue_SkyFlatScalings.ob --- cycles through all blue filters while keeping red filter set at V-BESSEL
    • Blue_SkyFlatScalings2.ob --- repeats the last 3 blue filters, since the filter changes in Blue may mean that the last few filters in Blue are done after and are non-simultaneous with Red.
    • Red_SkyFlatScalings.ob --- cycles through all red filters while keeping blue filter set at V-BESSEL
    • Red_SkyFlatScalings2.ob --- repeats the last few red filters
  • It would also be helpful to run the OBs generated by /home/LBTO/Calib_OBs/Mkskyflat1/mkskyflat1, with no exposure times specified on the command line, i.e. ./mkskyflat1.pl [e|m] 0 Bfilter Rfilter in home/LBTO/Calib_OBs/Mkskyflat1/ to test the exposure time scalings in that script.

Pointing Check
  • OBs in /home/LBTO/RB_VV_PointingOBs/
  • Slew to pointing star from list
  • Use LBTtools.Observe.lbcrangebal to correct pointing and co-pointing. If monocular, use LBTtools.LBC.ptautoadjust to correct pointing.

Measure focal plane tilt.
  • Slew to a field with relatively many stars and at reasonable elevation, e.g. elevation~75 degrees.
  • Obtain focus sequences for at least 2 position angles (0 and 180, e.g.)
    • OBs for this are RB_rVsuperfoc.ob and RB_rVsuperfoc180.ob in /home/LBTO/Calib_OBs/SUPERFOC/
  • Analyze in two ways:
    • run nmisc.starfoc for stars in subregions of each chip, usually top, middle, bottom, to produce an array of 12 best-focus values.
    • use focfit3.cl script in LBTtools.Sandbox and gnuplot to generate and display this array of best-focus values. You must load the IRAF package LBC as well as Sandbox, even though this task is in Sandbox.
      • (focfit3 uses Source Extractor to measure average FWHM in a bin. The number of bins along X and Y can be set).

Test existing and new collimation models.
  • Initialize LBC collimation models (done implicitly). Existing models are SXPMLBCCollimation.dat and DXPMLBCCollimation.dat (copied also to SXPMLBCCollimation.20110904.dat and DXPMLBCCollimation.20110904.dat).
  • Slew to a field. Collimate. Note Active Optics offsets needed to achieve collimation (i.e. how much correction was made to the model?)
  • Slew to zenith.
  • Repeat first two steps for about 10 fields separated by 10-15 degrees in elevation.
  • Next load the new collimation models. These are currently in /home/okuhn/. cp /home/okuhn/SXPMLBCCollimation.dat /home/telescope/TCS/Configuration/PSF/ and same for DX.
  • Click Initialize on the Left and Right PSF GUIs (this must be done, since we overwrote the existing models).
  • Slew to a field. Collimate. Note Active Optics offsets needed to achieve collimation.
  • Slew to zenith
  • Repeat first two steps for about 10 fields separated by 10-15 degrees in elevation.

Collect data to make co-pointed collimation models.
  • Slew to pointing star. Note AO offsets and total collimation
  • Run lbcrangebal
  • Collimate. Note AO offsets and total collimation.
  • Slew to another pointing star. Try to find stars at a range of elevations from 20-90 in steps of about 10 degrees.

Measure Rotator Center
  • Slew to a star field at moderate elevation and collimate
  • Take a pair of 5-sec images at two diametrically opposed position angles (0 and 180, e.g.). Good idea to obtain at least two pairs.
    • Ask the operator to slew.
    • OBs are rotatorcenter_pa0.ob and rotatorcenter_pa180.ob (these do not slew the telescope) in /home/LBTO/Calib_OBs/ROTCEN_Check/
  • Run script LBTtools.LBC.rotcenter to compute rotator centers.
    • Update eight header template files, blue_ehu_sci{1,2,3,4}.hdr and red_ehu_sci{1,2,3,4}.hdr on CMU: /home/lbccontrol/conf/ by updating the values for CRPIX1 and CRPIX2 in lines 2 and 3.
    • Define:
      • Delta_X = RotCen_X(new) - RotCen_X(old) and
      • Delta_Y = RotCen_Y(new) - RotCen_Y(old)
      • Then for chips 1,2,3:
        • CRPIX1(new) = CRPIX1(old) + Delta_X and
        • CRPIX2(new) = CRPIX2(old) + Delta_Y
      • And for chip 4:
        • CRPIX1(new) = CRPIX1(old) + Delta_Y and
        • CRPIX2(new) = CRPIX2(old) - Delta_X

Collect data for Pointing Model
A 28 star pointing run would be good for a "spring" set of pointing data. We have winter sets and can investigate any temperature dependence of the pointing model
  • Insure Rotator Center is correct in headers (CRPIX1, CRPIX2 of chip 2 are the most recent values for the rotator center)
  • Make sure the correct, most recent, collimation model is loaded.

Measure filter focus offsets
These should not change with the hub move, but it would be helpful to take some focus sequences.
  • OBs in /home/LBTO/Calib_OBs/SUPERFOC/
  • Make image list (in real time) with LBTtools.LBC.foclist
  • Run nmisc.starfoc to measure best focus for the sequence. Concentrate on region around rotator center. Be consistent.
  • Procedure for obtaining these data:
    • Slew to a field with many stars, though not too dense
    • Run dofpia
    • Run a few superfoc sequences using different filter pairs but always include the reference pair using the rV filters (RB_rV_superfoc.ob)
    • Repeat these two steps (dofpia, superfoc sequences).
      • Be sure to repeat each filter 3 or more times; and
      • Be sure to always include a reference pair using the r-SLOAN and V-BESSEL filters.

Test guiding, dithering (need to have a quick reduce script)
  • Develop OB - plan to have guide/focus stars on tech chips for all dither positions.
  • Check/correct pointing/co-pointing before: lbcrangebal script
  • SN target at 16hrs
  • M16 at 18hrs
  • There are additional targets in Calib_OBs/ASTRONtest

Check or refine calibrations if necessary:
  1. Astrometry - obtain one astrometric field. OBs are in /home/lbcobs/REDCOMMISSIONING/ASTROMETRIC_STD/
  2. Throughput - obtain one standard star field if photometric. OBs are in Calib_OBs package, in STANDARDS directory.

Investigate recently uncovered or ongoing problems with LBC or telescope:
Implement tools to aid observers and LBTO staff

Observing Plan (copy and paste this to the nightly logs)

  • Closed dome SW tests, if requested
  • Open dome SW tests, if requested

  • Run a "science" OB which dithers, guides (M16)
    • Create OB to put guide stars on both tech chips today.
    • Before running OB
      • Slew to nearby pointing star and correct pointing IE/CA
      • Correct copointing with Red and Blue images of pointing star
      • Load science OB and slew to target position
      • collimate/focus
      • Run science OB

  • Check existing collimation model
    • Take new data if necessary
    • Stable conditions!

  • Check existing pointing model
    • Implement new collimation model if necessary
    • Stable conditions!
    • Take 12 points, with pointing log on...
    • Take full set (56 points) if the first 12 points are off.

-- OlgaKuhn - 27 May 2010

-- OlgaKuhn - 02 Sep 2011

  • Red and Blue HK temperatures (thermal link) vs detector temperatures 29-31 August 2011:
    comparehkdet.jpg

  • HK Temperatures 27 June 2011:
    hktemp_27jun11.jpg

  • HK Temperatures 4 July 2011:
    hktemp_4jul11.jpg
I Attachment Action Size Date Who Comment
comparehkdet.jpgjpg comparehkdet.jpg manage 73 K 02 Sep 2011 - 23:08 OlgaKuhn Red and Blue HK temperatures (thermal link) vs detector temperatures 29-31 August 2011
hktemp_27jun11.jpgjpg hktemp_27jun11.jpg manage 71 K 02 Sep 2011 - 23:26 OlgaKuhn HK Temperatures 27 June 2011
hktemp_4jul11.jpgjpg hktemp_4jul11.jpg manage 69 K 02 Sep 2011 - 23:27 OlgaKuhn HK Temperatures 4 July 2011
Topic revision: r4 - 04 Sep 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback